# Beyond Walking: A Large-Scale Image-Text Benchmark for Text-based Person Anomaly Search

### v1

@@ 6 Conclusion
This work introduces the task of Text-based Person Anomaly Search, aimed at addressing the limitations of traditional text-based person retrieval in identifying abnormal behaviors. Given the lack of a benchmark in this domain, we construct a large-scale Pedestrian Anomaly Behavior (PAB) benchmark. The test data is sourced from real-world videos for practical evaluation, while the large-scale training set is generated using a diffusion model. We propose a cross-modal pose-aware framework that effectively integrates human pose patterns with identity-based hard negative pair mining to distinguish between normal and anomalous actions. Extensive experiments on the PAB benchmark verify the effectiveness of our Cross-Modal Pose (CMP) model in retrieving abnormal behaviors in real-world scenarios.
@@ Appendix A More Benchmark Details
A.3 Visualizations.
@@ Appendix B Experiment Details and Further Experiments
Method R@1 R@5 R@10 mAP
RaSa [5] 49.34 81.24 88.67 63.57
CMP (Ours) 72.80 96.01 97.47 83.47
Table 4: Comparisons with state-of-the-art text-based person search method in the text-based person anomaly search evaluation setting.
B.3 More Comparisons with state-of-the-art text-based person search methods.
In Table 4, we compare our method with the state-of-the-art text-based person search method, RaSa [5]. After 30 epochs of training on the same $0.1M$ PAB data, our method outperforms RaSa [5] by $23.46\%$ R@1 and $19.90\%$ mAP.
B.4 More Qualitative Result Examples.

### v2

@@ 6 Conclusion
We introduce Text-based Person Anomaly Search, a new task addressing the limitation of traditional retrieval in identifying anomalies. To support this, we construct the Pedestrian Anomaly Behavior (PAB) benchmark, combining real-world test videos with synthetic training data. PAB provides a large-scale training set with fine-grained annotations and high anomaly ratio, facilitating the model learning. Our Cross-Modal Pose-aware (CMP) framework integrates pose patterns and hard negative mining to distinguish between normal and anomalous actions. Extensive experiments on PAB and UCC verify the effectiveness of our Cross-Modal Pose (CMP) model in retrieving both anomalous and normal behaviors in real-world scenarios.
Supplementary Material
Figure 6: Dataset Statistics. An overview of the attribute annotations, including the distribution of categories across the training and test sets. Specifically, it covers normal action categories (a, d), anomaly categories (b, e), and scene categories (c, f). Due to the natural long-tail distribution of the data and the space limitation, we present the top 15 most common classes for each category to ensure clarity. (Best viewed when zooming in.)
Datasets Modality #Frames #Scenes #Anomaly Types Anomaly:Normal Avg Word Open Set Data Source
UCSD Ped2 [26] Video 4,560 1 5 Classes 1:2 - $\checkmark$ Collection
UMN [34] Video 7,741 3 1 Classes 1:4 - $\checkmark$ Collection
UCSD Ped1 [26] Video 14,000 1 5 Classes 1:2 - $\checkmark$ Collection
CUHK Avenue [31] Video 30,652 1 5 Classes 1:7 - $\checkmark$ Collection
Subway Exit [2] Video 64,901 1 3 Classes 1:13 - $\checkmark$ Collection
Subway Entrance [2] Video 144,250 1 5 Classes 1:11 - $\checkmark$ Collection
Street Scene [38] Video 203,257 1 17 Classes 1:4 - $\checkmark$ Collection
UBnormal [1] Video 236,902 29 22 Classes 2:3 - $\checkmark$ Synthesis
ShanghaiTech [32] Video 317,398 13 11 Classes 1:18 - $\checkmark$ Collection
UCF-Crime [43] Video 13,741,393 Unlimited 13 Classes $\ll$1:1 - $\times$ Collection
UCA [53] Video, Text 13,741,393 Unlimited 13 Classes $\ll$1:1 20.2 $\times$ Collection
PAB (Ours) Image, Text 1,015,583 480 1600 Classes 3:2 52.0 $\checkmark$ Synthesis &
Collection
Table 6: Comparison of the statistics of our PAB and other Video Anomaly Detection (VAD) datasets. The statistics of previous datasets have been recorded in [1].
Figure 7: Dataset Examples. 12 training (synthetic) image-text pairs from the PAB dataset are at the top, while 12 test (real-world) image-text pairs are at the bottom. Half of the examples depict anomaly behaviors, while the other half show corresponding normal actions. Each pair is annotated with scene and action (or anomaly) classes. Minor errors may be present in the generated captions of the training set, whereas the captions in the test set have been refined by professionals. (Best viewed on a computer screen with zoom.)
@@ Appendix A More Benchmark Details
A.3 Comparisons with More Video Anomaly Detection Datasets.
In Table 6, we compare our proposed PAB dataset with the most utilized Video Anomaly Detection (VAD) datasets. Eight metrics are reported: modality, number of frames/images, scenes, anomaly types, the proportion of anomaly versus normal, average word of each text, open-set characteristics, and data source. Compared to other video datasets, PAB is distinguished as an image-text pair dataset and features a higher number of anomaly types from a broader range of event scenes. For all datasets, the ”Anomaly:Normal” ratio represents the proportion of anomaly video frames/images to normal video frames/images. While most VAD datasets are annotated solely with normal/abnormal labels or abnormal category labels, PAB provides detailed annotations including appearance descriptions, actions, and scene information. Most video datasets maintain open-set characteristics for anomaly detection. To ensure consistent open-set characteristics, we provide a real-world Out-of-Distribution (OOD) test set for PAB sourced from UCF-Crime [43]. Notably, UBnormal [1] is also a synthetic dataset, but unlike PAB, both its training and test sets consist entirely of synthesized data.
A.4 Visualizations.
@@ Appendix B Experiment Details and Further Experiments
B.3 More Qualitative Result Examples.